OpenStudy (anonymous):

Two identical point charges exert an attractive force of 0.0010N on each other when they are 0.60m apart. Can someone please help me calculate the magnitude of each charge?

OpenStudy (farcher):

Use Coulomb's inverse square law equation.

OpenStudy (anonymous):

@Farcher is that F=(kq(1)q(2))/r^2?

OpenStudy (michele_laino):

please you have to use the Coulomb's law: \[F = \frac{1}{{4\pi {\varepsilon _0}}}\frac{{{q^2}}}{{{d^2}}}\]

OpenStudy (michele_laino):

\[\frac{1}{{4\pi {\varepsilon _0}}} = 9 \cdot {10^9}\quad \frac{{N{m^2}}}{{{C^2}}}\]

OpenStudy (michele_laino):

what is q?